Just after opening a parachute of negligible mass, a parachutist of mass 91.0 kg experiences an instantaneous upward acceleration of 0.87 m/s2. Find the force of the air on the parachute.

magnitude:

direction:
Answer #1

Solution:

Let the upward force of the air = Fa

(Fa > weight (mg) as there is an upward acceleration [= deceleration downward])

The resultant force on the body Fr = (Fa - mg) ↑

Fr = ma↑ = (Fa - mg)↑

Fa = m(a + g)↑ .. (ie Fa must = the weight [mg] and provide the accelerating force [ma])

Fa = 91.0 (9.81 + 0.87) ↑

Fa = 971.88 N upward

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions.